OpenStudy (some.random.cool.kid):

Which exponential function goes through the points (1, 8) and (4, 64)? f(x) = 4(2)x f(x) = 2(4)x f(x) = 4(2)-x f(x) = 2(4)-x

OpenStudy (some.random.cool.kid):

there

OpenStudy (danjs):

alrighty

OpenStudy (some.random.cool.kid):

:}

OpenStudy (danjs):

An exponential function has a general form of Y = C*a^x

OpenStudy (some.random.cool.kid):

yes

OpenStudy (danjs):

\[y = C*a^x\]

OpenStudy (danjs):

points given are (1,8) and (4,64) right?

OpenStudy (some.random.cool.kid):

yes

OpenStudy (danjs):

so since we have 2 points, we can form 2 equations from that general form.. \[8 = C*a^1 ~~~~~and~~~~~64 = C*a^4\]

OpenStudy (danjs):

see that?

OpenStudy (some.random.cool.kid):

yes ok, so what should i do next.

OpenStudy (danjs):

now we have 2 equations with 2 unknowns, a and C, we can solve them and figure out a and C!

OpenStudy (danjs):

oops i got excited exclamation. lol

OpenStudy (some.random.cool.kid):

:} ok

OpenStudy (danjs):

Lets take the first equation and solve it for C \[8 = C*a^1\]

OpenStudy (danjs):

\[C = \frac{ 8 }{ a }\]

OpenStudy (danjs):

you get that part?

OpenStudy (some.random.cool.kid):

alright so far im still with you ;}

OpenStudy (danjs):

ok, since we have C=8/a , lets put that value into the second equation... \[64 = C*a^4\] \[64 = \frac{ 8 }{ a }*a^4\]

OpenStudy (danjs):

good with that?

OpenStudy (some.random.cool.kid):

ok

OpenStudy (some.random.cool.kid):

now we solve a right

OpenStudy (danjs):

exactly, what you have so far is \[64 = 8*\frac{ a^4 }{ a }\] divide both sides by 8 first

OpenStudy (some.random.cool.kid):

speaking of which i did 8 x 8 because i was trying to plug in my choices, and for my answer i chose A because i was wondering when talking about the ^x i figured i would do what you displayed above.

OpenStudy (danjs):

I see, you could plug into the answers and see what works, but you should learn how to do this... what if no multiple choice is given? .lol we are almost there i promis so after dividing by 8 on both sides you get \[8 = \frac{ a^4 }{ a }\]

OpenStudy (danjs):

\[\frac{ a^4 }{ a } = \frac{ a^4*a ^{-1} }{ 1 } = a ^{4-1} = a ^{3}\]

OpenStudy (some.random.cool.kid):

oh i wasnt worried about rushing im in no hurry i was just wondering if that tactic would work but yeah your method works for me as well.

OpenStudy (danjs):

a^3 = 8 the cubed root of 8 is 2..... 2*2*2 = 8 so we figured out that a = 2

OpenStudy (danjs):

you see that part?

OpenStudy (some.random.cool.kid):

yes

OpenStudy (some.random.cool.kid):

i figured that part

OpenStudy (danjs):

now we have 'a' y = C*a^x y = C*2^x we can find out what C is now

OpenStudy (danjs):

What were the original 2 points given?

OpenStudy (danjs):

(1,8) and (4,64)

OpenStudy (some.random.cool.kid):

(1, 8) and (4, 64)

OpenStudy (danjs):

we can use one of those points in the equation, and figure out C use ( 1 , 8) the power of a^1 is easier than a^4 from the other point

OpenStudy (some.random.cool.kid):

tjhats kind of what i did.

OpenStudy (some.random.cool.kid):

thats*

OpenStudy (danjs):

y = C*2^x point(x,y) = (1,8) 8 = C*2^1 8 = 2C C = 4

OpenStudy (danjs):

overall equation , now that we found 'a' and C is \[\large\color{blue}{ y = 4*2^x}\]

OpenStudy (some.random.cool.kid):

yes i solved that too however the putting it all together is what stumbled me. but yeah now i see it.

OpenStudy (some.random.cool.kid):

so was i correct for picking A originally.

OpenStudy (some.random.cool.kid):

if so then im :}

OpenStudy (danjs):

cool, so now anytime you get a question that gives you 2 points and says it is exponential (y=C*a^x) , you can figure out the equation! right?

OpenStudy (danjs):

1) Make an equation using each point 2) solve one equation for C 3) put that C into the second equation 3) Solve that for a = 4) Now you have a, use one point and that 'a' in the y=C*a^x to find C 5) done

OpenStudy (danjs):

Go back through and right out all the steps on paper, it will be clear

OpenStudy (some.random.cool.kid):

done i wrote it

OpenStudy (danjs):

cool, if you want to run through another for practice, ill help out if you want

OpenStudy (some.random.cool.kid):

ok sure i guess i have free time.

OpenStudy (danjs):

k

OpenStudy (some.random.cool.kid):

did you have one in mind or should i pull one up.

OpenStudy (some.random.cool.kid):

up to you :}

OpenStudy (triciaal):

@DanJS thanks

OpenStudy (some.random.cool.kid):

change of plan got to go maybe another day? bye and thanks again! :}

OpenStudy (danjs):

@triciaal you are welcome, what are you thanking me for? lol

OpenStudy (triciaal):

@DanJS responding when I tagged patiently helping the "cool kid" reminding me about the function notation y = c*a^x that the poster meant.